OpenStudy (anonymous):

write a function in slope-intercept form whose graph satisfy the given equation... slope = -2, passing through (-4,3)

OpenStudy (anonymous):

y=-2x+3

OpenStudy (saifoo.khan):

y = mx + b 3 = -2(-4) + b b = 3 -8 b = -5 y = mx + b y = -2x -5 <---- answer.